40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 21 17:01:59 2022, Paco, Update, BHD, BH55 RFPD installed - part I IMG_3760.jpegIMG_3764.jpeg
    Reply  Thu Sep 22 19:51:58 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part II 
       Reply  Fri Sep 23 19:04:12 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part III 
          Reply  Thu Sep 29 18:01:14 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part IV 
             Reply  Fri Sep 30 18:30:12 2022, Anchal, Update, ASS, Model Changes 
                Reply  Wed Oct 12 14:01:23 2022, Radhika, Update, ASS, Model Changes YARM_ASS.png
                   Reply  Wed Oct 12 18:18:16 2022, rana, Update, ASS, Model Changes 
    Reply  Fri Sep 23 14:10:19 2022, Radhika, Update, BHD, BH55 RFPD installed - part I IMG_3766.jpegIMG_3770.jpegIMG_3773.jpeg
       Reply  Fri Sep 23 18:31:46 2022, rana, Update, BHD, BH55 RFPD installed - part I 
          Reply  Mon Sep 26 11:39:37 2022, Paco, Update, BHD, BH55 RFPD installed - part II PXL_20220926_175010061.MOTION.jpgBH55_Transimpedance_Measurement.pdf
             Reply  Thu Oct 6 11:12:14 2022, Anchal, Update, BHD, BH55 RFPD installation complete 
Message ID: 17157     Entry time: Fri Sep 23 19:04:12 2022     In reply to: 17152     Reply to this: 17165
Author: Anchal 
Type: Update 
Category: BHD 
Subject: BH55 LSC Model Updates - part III 

BH55

I further updated LSC model today with following changes:

  • BH55 whitening switch binary output signal is now routed to correct place.
    • Switching FM1 which carries dewhitening digital filter will always switch on corresponding analog whitening before ADC input.
  • The whitening can be triggered using LSC trigger matrix as well.
  • The ADC_0 input to LSC subsystem is now a single input and channels are separated inside the subsystem.

The model built and installed with no issues.

Further, the slow epics channels for BH55 anti-aliasing switch and whitening switch were added in /cvs/cds/caltech/target/c1iscaux/C1_ISC-AUX_LSCPDs.db


IPC issue resolved

The IPC issue that we were facing earlier is resolved now. The BH55_I and BH55_Q signal after phase rotation is successfully reaching c1hpc model where it can be used to lock LO phase. To resolve this issue, I had to restart all the models. I also power cycled the LSC I/O chassis during this restart as Tega suspected that such a power cycle is required while adding new dolphin channels. But there is no way to find out if that was required or not. Good news is that with the new cds upgrade, restarting rtcds models will be much easier and modular.


ETMY Watchdog Updated

[Anchal, Tega]

Since ETMY does not use HV coil driver anymore, the watchdog on ETMY needs to be similar to other new optics. We made these updated today. Now ETMY watchdog while slowly ramps down the alignment offsets when it is tripped.

ELOG V3.1.3-